# Physical Correlations and Predictions Emerging from Modern Core-Collapse Supernova Theory

### 12 Jan 2024 v1

@@ 1 Introduction
Building on the early pioneering work of Colgate & White (1966), Arnett (1967), Wilson (1971), Bowers & Wilson (1982), and Bethe & Wilson (1985), with the recognition of the importance of neutrino-driven turbulence and convection (Herant et al., 1994; Burrows et al., 1995; Janka & Mueller, 1996), and paralleling the developments in neutrino physics and supercomputers over the decades (Burrows, 2019), core-collapse supernova (CCSN) theory has emerged after 60 years of progress to become a mature, if complicated and multi-physics, international enterprise. Neutrino heating, aided by the stress of neutrino-driven turbulence, is accepted as the key mechanism of explosion. However, turbulence is chaotic and this implies that a range of outcomes and observable values are likely, even for the same massive star progenitor. The explosive expansion of the initially stalled bounce shock wave seems to be a critical phenomenon (Burrows & Goshy, 1993) and is associated with the shock’s pulsational instability, but a robust analytic model has proven elusive. This has necessitated detailed multi-dimensional numerical simulations, whose complexity and expense had slowed conceptual and quantitative progress. The accretion before explosion of the density jump associated with the Si/O interface in the massive star progenitor, itself correlated with “compactness” (see §2.1), oftimes inaugurates explosion (Vartanyan et al., 2018; Wang et al., 2022). However, this is not always the case, and its invocation as an exclusive trigger for explosion should be nuanced. Curiously, a quick explosion does not in general lead to an energetic explosion (see §3.1).
@@
Interestingly, in recent late-time 3D simulations, Wang & Burrows (2023a) observe aspherical winds (Duncan et al., 1986; Burrows, 1987a; Burrows et al., 1995; Qian & Woosley, 1996) emerging from the PNS after a delay of seconds following the onset of explosion; these produce secondary shock waves upon encountering the primary ejecta. As much as 10% of the explosion energy can be due to these winds. However, other groups have yet to simulate to late enough times after bounce to verify the details of this phenomenon, itself also associated with the important nucleosynthetic phase of -rich freeze-out (Woosley et al., 2002; Nomoto et al., 2013; Arcones & Thielemann, 2023; Wang & Burrows, 2023b). Furthermore, which stars leave neutron stars and which black holes is still not resolved, though it is known that previous mappings were too simplistic (Burrows, 1987b; Heger et al., 2003; Sukhbold et al., 2016). There are new ideas on this partitioning (Burrows et al., 2023a), but these have not been thoroughly vetted and depend a great deal on the still-evolving progenitor models inherited.
Despite these many remaining issues and the physical and numerical complexity of the phenomenon, and despite the fact that chaotic turbulence renders a clear one-to-one mapping of progenitor to outcome naturally blurred, there has emerged a broad consensus on the viability of the turbulence-aided neutrino-driven mechanism of core-collapse supernovae (CCSNe). Modern sophisticated 3D simulations witness neutrino-driven explosions naturally and without artifice (Lentz et al., 2015; Roberts et al., 2016; Takiwaki et al., 2016; Müller et al., 2017; Ott et al., 2018; Glas et al., 2019; Müller et al., 2019; Burrows et al., 2019, 2020; Müller & Varma, 2020; Powell & Müller, 2020; Stockinger et al., 2020; Kuroda et al., 2020; Bollig et al., 2021; Vartanyan et al., 2022, 2023) and numerous reviews have articulated the mechanism’s details (Janka, 2012; Burrows, 2013; Burrows et al., 2018, 2020; Burrows & Vartanyan, 2021; Wang et al., 2022). With this paper, we summarize the physical and observable correlations that have emerged from the panoramic view enabled by our new and unprecedented set of twenty state-of-the-art 3D core-collapse simulations taken to late times.
@@ 2 Initial Progenitor Structures
For this study we employ solar-metallicity, initially non-rotating, spherical massive star progenitor models from Sukhbold et al. (2016) and Sukhbold et al. (2018) and we do not follow magnetic fields. The Fornax code and calculational details are described in the Appendix and in Skinner et al. (2019). As discussed in Wang et al. (2022), it is predominantly the density profile at collapse that determines explodability and this profile is not rigorously monotonic with zero-age main-sequence (ZAMS) mass. Figure 1 depicts these mass density profiles for the models we include in this paper. As one sees, while the steepness of the envelope in the Chandrasekhar-like core is roughly an increasing function of ZAMS mass, it is not rigorously so, nor is the “compactness” parameter (O’Connor & Ott, 2011) (see §2.1); we will find in this paper a tighter relation of the observables with compactness than with ZAMS mass (§5).
Importantly, there remain many ambiguities in the initial progenitor models and, as important as they are to core-collapse supernova theory, the progenitor models can not yet be considered definitive. Issues such as the C(,) rate, binarity, overshoot, wind mass loss, shell mixing, semi-convection, magnetic field effects, and rotation are still to be retired. Moreover, a new generation of 3D progenitor models (Couch et al., 2015; Jones et al., 2016; Chatzopoulos et al., 2016; Müller et al., 2016; Jones et al., 2017; Yoshida et al., 2019; Fields & Couch, 2020; Bollig et al., 2021; McNeill & Müller, 2021; Fields & Couch, 2021; Varma & Müller, 2021; Yoshida et al., 2021) is emerging. As a result of these many remaining questions and uncertainties, models for the stellar evolution of massive stars to the point of collapse can not be considered to have converged. Nevertheless, we use the Sukhbold et al. (2016) and Sukhbold et al. (2018) model suite as the best and most comprehensive available. Moreover, we suggest that our derived correlations between observables and compactness may be usefully robust, providing ongoing insight as the progenitor model studies evolve into the future.
@@ 2.1 Compactness
The shallowness of the mass density profiles depicted in Figure 1 is related to the post-bounce mass accretion rate history and the accretion component of the neutrino luminosities. It is through these quantities that the complicated dynamics of the supernova mainly rests. One crude, but useful, metric of this shallowness is the compactness. Compactness (O’Connor & Ott, 2011)222Introduced to explore black hole formation, it is defined as , for which we prefer to set equal to 1.75 when studying supernova explosions, not , as did O’Connor & Ott (2011). is not a predictor of explodability (Wang et al., 2022). Nevertheless, it is a useful, if gross, measure of progenitor structure and correlates well with numerous observables and physical characteristics of supernova dynamics.
@@ 3 Basic Explosion Results
To set the stage for the subsequent comparisons, we display in Figure 4 the temporal evolution of the averaged shock radius (top left), PNS baryon mass (top right), explosion energy (bottom left), and kick speed (bottom right). Note that explosion occurs over a range of post-bounce times from 100 milliseconds (ms) to 800 ms, with the lower mass progenitors generally exploding more quickly and the more massive progenitors taking longer. In fact, the shock waves for the 9 to 9.5 progenitors barely stall, while that for the 20 progenitor recedes significantly before exploding. We note that among those models for which the shock initially recedes it is only those for which the compactness and associated mass accretion rate and accretion luminosity are high that explode. The 12.25 and 14 models have too low a compactness to reverse the recession of their stalled accretion shocks. As Figure 4 indicates, the baryon mass asymptotes earlier than the energy and kick speeds. The latter both take many seconds to “flatten.” The explosion energy (here including the stellar mantle overburden binding energy) starts negative and can take more than a second to become positive. Most of the neutrino energy deposited early on by absorption goes into overcoming the gravitational term and only later is the positive value necessary for explosion achieved. Those numerical simulations in papers that stop within 1 second of bounce can not possibly provide an informed estimate of the explosion energy. The same can be said for the kick speed (Coleman & Burrows, 2022; Burrows et al., 2023b) as Figure 4 reveals, models halted within a second of bounce can not provide a useful final kick magnitude.
@@ 3.1 Time of Explosion
Figure 5 displays the relation between the compactness parameter and the explosion time. We see that the relation between the two is very roughly monotonic models with larger compactness take longer to explode, but there is a lot of scatter. This scatter may in part reflect the chaos of the dynamics, but the degree to which this is true has yet to be determined. We emphasize that this theme of chaos, stochasticity, and scatter in the correlations suffuses our results and may reflect the very real scatter in the outcomes, even for the same progenitor. In fact, one expects a distribution function in the observables for the same progenitor, arising from the turbulent chaos in the dynamics. The width and form of such distribution functions has yet to be determined and is an important topic for future investigation. One of the motivations for this study of many models, as opposed to only a few, is the idea that in so doing one can glimpse the various correlations that would otherwise be obscured. We think we have achieved this in this paper, but clearly an even more expanded (and expensive!) investigation would be better.
@@ 4 Explosion and Debris Morphology
The aspherical morphology of the explosion debris reflects the turbulence of the pre- and post-explosion dynamics that has been a central feature of CCSN theory for decades (Burrows et al., 1995)444The Standing Accretion Shock Instability (SASI) (Blondin et al., 2003), a vortical-acoustic instability (Foglizzo et al., 2007), once seemed manifest in most CCSN simulations. However, the associated up and down “sloshing” was an artifact of the axial effect in 2D simulations; the turbulence witnessed was in fact neutrino-driven convection. However, Foglizzo et al. (2006) has shown that when the stalled shock recedes a SASI can emerge, but current 3D modeling (Burrows et al., 2020, 2023a) reveals it likely to be the spiral SASI (Blondin & Shaw, 2007), instead of the classic SASI. Its emergence can presage a late-time explosion (Andresen et al., 2019), an explosion followed by the formation of a black hole (Burrows et al., 2023a), or (as we see quite often (Burrows et al., 2023a)), a fizzled collapse that will over time lead to the “quiescent” collapse to a black hole.. Bubble structures are universally in evidence (see Sato et al. (2021)), with larger bubbles generally driving the explosions. Importantly, the longer the delay to explosion, the more vigorous the turbulence achieved, with the Mach number of the convective overturning motions prior to explosion getting as high as 0.5. However, if the explosion is more prompt, the post-shock turbulence does not have sufficient time to grow to such vigor and large convective structures are not produced. The result in such cases is a more spherical explosion, albeit with many small bubble structures. Given the trends of compactness and progenitor ZAMS mass with explosion time exposed in Figure 5, we would expect lower compactness, lower ZAMS mass progenitors would explode more spherically, while those with larger compactness would explode more aspherically.
@@
Such structures have a direct bearing on the relative line profiles and the spatial distribution of the elements in supernova debris (Taubenberger et al., 2009; Sato et al., 2021; Fang et al., 2023) and the associated (anti)correlation of Ni and progenitor oxygen shell material is a qualitative prediction of these 3D models. Moreover, one would expect a correlation with the polarization of the inner debris (Leonard et al., 2000, 2006; Tanaka et al., 2017; Nagao et al., 2023). Figure 9 portrays the relationship between the explosion energy and the dipole anisotropies of the ejecta mass (left) and the ejecta energies (right) among our set of twenty models. These panels suggest a roughly monotonic correlation of two indices of anisotropy with explosion energy that is a new prediction that emerges from our study. We expect that, though with the common caveat of some scatter, more energetic explosions to evince greater anisotropy, but this needs to be tested.
@@ 5.2 Explosion Energy
Two of the most important observables of CCSN theory are the explosion energy and the gravitational mass of the residual neutron star. Our theoretical results reveal that the two are tightly coupled. Figure 11 portrays the relation between the two and indicates that, with the caveat that some of our models have not quite asymptoted to their final explosion energies (but are generally within 10% or better), the explosion energy is a strictly monotonic function of neutron star gravitational mass. This relationship between two observables is one of the central results of this paper and speaks to the importance of our philosophy to explore many long-term 3D state-of-the-art simulations to extract meaningful predictions. This strong mapping between observables would not have been clearly discerned without a 3D investigation of many progenitors to late times and should inform future observational campaigns to extract physically useful parameters of core-collapse supernova explosions. The study of but one single model would not be so illuminating.
Figure 12 plots the connection between binding energy and explosion energy (left) and compactness and explosion energy (right). Both relationships are roughly monotonic and robust, but the explosion energy/binding energy correlation is physically one of the most important findings from the theoretical perspective. It implies that the supernova energy is a “self-regulating” phenomenon, akin in broad outline to winds. Such self-regulation, mediated through the shared implicit dependence upon compactness, may become one of the unifying principles of CCSN theory going forward.
@@ 5.3 Ni Yields and Nucleosynthesis
Supernova nucleosynthesis is a vast subject that spans almost eight decades. Components are pre-supernova progenitor yield calculations, explosive nucleosynthesis, -rich freeze-out, and various sub-processes of differing import. Reviews can be found in Woosley et al. (2002); Nomoto et al. (2013); Arcones & Thielemann (2023). The decades-long tradition of 1D models Thielemann et al. (1996); Rauscher et al. (2002); Woosley et al. (2002); Heger & Woosley (2010); Nomoto et al. (2013); Sukhbold et al. (2016); Limongi & Chieffi (2018); Curtis et al. (2019) is now being challenged by 3D models (Lentz et al., 2015; Wongwathanarat et al., 2015; Harris et al., 2017; Wanajo et al., 2018; Sieverding et al., 2020, 2023; Wang & Burrows, 2023a, b). Winds play an important role (Duncan et al., 1986; Burrows, 1987a; Burrows et al., 1995; Qian & Woosley, 1996; Wang & Burrows, 2023a), as do neutrinos (Woosley et al., 1990; Fröhlich et al., 2006; Pruet et al., 2006) and -ray emissions of the ejecta can play a diagnostic role (Diehl & Timmes, 1998; Diehl et al., 2021; Diehl & Prantzos, 2023).
One of the central observables of core-collapse supernova explosions is the Ni yield (in solar masses), since the Ni decay chain to Fe powers some of the supernova light curve and can easily be measured. Hence, the predicted Ni yields are central physical quantities of the CCSN phenomenon. Figure 13 provides the relation that emerges in our 3D simulation suite between explosion energy (in Bethes 10 ergs) and Ni production (in ), the latter derived using the formalism described in Wang & Burrows (2023b). This too is one of the central results of this paper, but the correlation between the two has already been anticipated theoretically (e.g. Sukhbold et al. (2016)) and observed astronomically (Rodríguez et al., 2023). The left hand side provides the Ni correlation with explosion energy, with the range of predicted energies due to the fact that not all 3D models have yet asymptoted to their final values indicated. The right hand side shows the same numerical data, but superposed on the Rodríguez et al. (2023) measurements. The overlap between the two suggests a reasonable concordance, though both trends are subject to natural variation. Note that the strong theoretical trend of Ni mass with explosion energy includes the exploding black hole formers (triangles).
@@
Figure 16 shows the dependence of the freeze-out mass upon compactness (top left), ejecta mass dipole (top right), and explosion energy (bottom). We see here too the rough monotonicity, in particular versus the explosion energy, of the mass in the freeze-out component. This follows the corresponding correlation of the explosion energy with Ni mass depicted in Figure 13 and is not surprising. It is nevertheless a strong prediction of the theory of CCSN explosions, as we see it, and is testable.

### 12 Mar 2024 v2

@@ 1 Introduction
Building on the early pioneering work of Colgate & White (1966), Arnett (1967), Wilson (1971), Bowers & Wilson (1982), and Bethe & Wilson (1985), with the recognition of the importance of neutrino-driven turbulence and convection (Herant et al., 1994; Burrows et al., 1995; Janka & Mueller, 1996), and paralleling the developments in neutrino physics and supercomputers over the decades (Burrows, 2019), core-collapse supernova (CCSN) theory has emerged after 60 years of progress to become a mature, if complicated and multi-physics, international enterprise. Neutrino heating, aided by the stress of neutrino-driven turbulence, is commonly accepted as the key mechanism of explosion. However, turbulence is chaotic and this implies that a range of outcomes and observable values are likely, even for the same massive star progenitor. The explosive expansion of the initially stalled bounce shock wave seems to be a critical phenomenon (Burrows & Goshy, 1993) and is associated with the shock’s pulsational instability, but a robust analytic model has proven elusive. This has necessitated detailed multi-dimensional numerical simulations, whose complexity and expense had slowed conceptual and quantitative progress. The accretion before explosion of the density jump associated with the Si/O interface in the massive star progenitor, itself correlated with “compactness” (see §2.1), oftimes inaugurates explosion (Vartanyan et al., 2018; Wang et al., 2022). However, this is not always the case, and its invocation as an exclusive trigger for explosion should be nuanced. Curiously, a quick explosion does not in general lead to an energetic explosion (see §3.1).
@@
Interestingly, in recent late-time 3D simulations, Wang & Burrows (2023a) observe aspherical winds (Duncan et al., 1986; Burrows, 1987a; Burrows et al., 1995; Qian & Woosley, 1996) emerging from the PNS after a delay of seconds following the onset of explosion; these produce secondary shock waves upon encountering the primary ejecta. As much as 10% of the explosion energy can be due to these winds. However, other groups have yet to simulate to late enough times after bounce to verify the details of this phenomenon, itself also associated with the important nucleosynthetic phase of -rich freeze-out (Woosley et al., 2002; Nomoto et al., 2013; Arcones & Thielemann, 2023; Wang & Burrows, 2023b). Furthermore, which stars leave neutron stars and which black holes is still not resolved, though it is known that previous mappings were too simplistic (Burrows, 1987b; Heger et al., 2003; Sukhbold et al., 2016; Ertl et al., 2016). There are new ideas on this partitioning (Burrows et al., 2023a), but these have not been thoroughly vetted and depend a great deal on the still-evolving progenitor models inherited.
Despite these many remaining issues and the physical and numerical complexity of the phenomenon, and despite the fact that chaotic turbulence renders a clear one-to-one mapping of progenitor to outcome naturally blurred, there has emerged a broad consensus on the viability of the turbulence-aided neutrino-driven mechanism of core-collapse supernovae (CCSNe). Modern sophisticated 3D simulations witness neutrino-driven explosions naturally and without artifice (Lentz et al., 2015; Roberts et al., 2016; Takiwaki et al., 2016; Müller et al., 2017; Ott et al., 2018; Glas et al., 2019; Müller et al., 2019; Burrows et al., 2019, 2020; Müller & Varma, 2020; Powell & Müller, 2020; Stockinger et al., 2020; Kuroda et al., 2020; Bollig et al., 2021; Vartanyan et al., 2022, 2023) and numerous reviews have articulated the mechanism’s details (Janka, 2012; Burrows, 2013; Müller et al., 2016a, 2017; Burrows et al., 2018, 2020; Mandel & Müller, 2020; Burrows & Vartanyan, 2021; Wang et al., 2022). With this paper, we summarize the physical and observable correlations that have emerged from the panoramic view enabled by our new and unprecedented set of twenty state-of-the-art 3D core-collapse simulations taken to late times.
@@ 2 Initial Progenitor Structures
For this study we employ solar-metallicity, initially non-rotating, spherical massive star progenitor models from Sukhbold et al. (2016) and Sukhbold et al. (2018) and we do not include magnetic fields. The Fornax code and calculational details are described in the Appendix and in Skinner et al. (2019). As discussed in Wang et al. (2022), it is predominantly the density profile at collapse that determines explodability and this profile is not rigorously monotonic with zero-age main-sequence (ZAMS) mass. Figure 1 depicts these mass density profiles for the models we include in this paper. As one sees, while the steepness of the envelope in the Chandrasekhar-like core is roughly an increasing function of ZAMS mass, it is not rigorously so, nor is the “compactness” parameter (O’Connor & Ott, 2011) (see §2.1); we will find in this paper a tighter relation of the observables with compactness than with ZAMS mass (§5).
Importantly, there remain many ambiguities in the initial progenitor models and, as important as they are to core-collapse supernova theory, the progenitor models can not yet be considered definitive. Issues such as the C(,) rate, binarity, overshoot, wind mass loss, shell mixing, semi-convection, magnetic field effects, and rotation are still to be retired. Moreover, a new generation of 3D progenitor models (Couch et al., 2015; Jones et al., 2016; Chatzopoulos et al., 2016; Müller et al., 2016b; Jones et al., 2017; Yoshida et al., 2019; Fields & Couch, 2020; Bollig et al., 2021; McNeill & Müller, 2021; Fields & Couch, 2021; Varma & Müller, 2021; Yoshida et al., 2021) is emerging. As a result of these many remaining questions and uncertainties, models for the stellar evolution of massive stars to the point of collapse can not be considered to have converged. Nevertheless, we use the Sukhbold et al. (2016) and Sukhbold et al. (2018) model suite as the best and most comprehensive available. Moreover, we suggest that our derived correlations between observables and compactness may be usefully robust, providing ongoing insight as the progenitor model studies evolve into the future.
@@ 2.1 Compactness
The shallowness of the mass density profiles depicted in Figure 1 is related to the post-bounce mass accretion rate history and the accretion component of the neutrino luminosities. It is through these quantities that the complicated dynamics of the supernova mainly rests. One crude, but useful, metric of this shallowness is the compactness. Compactness (O’Connor & Ott, 2011)222Introduced to explore black hole formation, it is defined as , for which we prefer to set equal to 1.75 when studying supernova explosions, not , as did O’Connor & Ott (2011). is not a predictor of explodability (Wang et al., 2022) in our simulations. Nevertheless, it is a useful, if gross, measure of progenitor structure and correlates well with numerous observables and physical characteristics of supernova dynamics.
@@ 3 Basic Explosion Results
To set the stage for the subsequent comparisons, we display in Figure 4 the temporal evolution of the averaged shock radius (top left), PNS baryon mass (top right), explosion energy (bottom left), and kick speed (bottom right). Note that explosion occurs over a range of post-bounce times from 100 milliseconds (ms) to 800 ms, with the lower mass progenitors generally exploding more quickly and the more massive progenitors taking longer. In fact, the shock waves for the 9 to 9.5 progenitors barely stall, while that for the 20 progenitor recedes significantly before exploding. We note that among those models for which the shock initially recedes it is only those for which the compactness and associated mass accretion rate and accretion luminosity are high that explode. The 12.25 and 14 models have too low a compactness to reverse the recession of their stalled accretion shocks. As Figure 4 indicates, the baryon mass asymptotes earlier than the energy and kick speeds. The latter both take many seconds to “flatten,” except in the case of the lowest-mass and lowest compactness progenitors. We note that explosion energies can ramp up quickly in some MHD explosions (Obergaulinger et al., 2018; Müller et al., 2018; Obergaulinger & Aloy, 2020). The explosion energy (here including the stellar mantle overburden binding energy) starts negative and can take more than a second to become positive. Most of the neutrino energy deposited early on by absorption goes into overcoming the gravitational term and only later is the positive value necessary for explosion achieved. Those numerical simulations in papers that stop within 1 second of bounce can not possibly provide an informed estimate of the explosion energy. The same can be said for the kick speed (Coleman & Burrows, 2022; Burrows et al., 2023b) as Figure 4 reveals, models halted within a second of bounce can not provide a useful final kick magnitude.
@@ 3.1 Time of Explosion
Figure 5 displays the relation between the compactness parameter and the explosion time. The explosion time in this work is defined as the time after which the shock never again recedes in radius. We see that the relation between the two is very roughly monotonic models with larger compactness take longer to explode, but there is a lot of scatter. This scatter may in part reflect the chaos of the dynamics, but the degree to which this is true has yet to be determined. We emphasize that this theme of chaos, stochasticity, and scatter in the correlations suffuses our results and may reflect the very real scatter in the outcomes, even for the same progenitor. In fact, one expects a distribution function in the observables for the same progenitor, arising from the turbulent chaos in the dynamics. The width and form of such distribution functions has yet to be determined and is an important topic for future investigation. One of the motivations for this study of many models, as opposed to only a few, is the idea that in so doing one can glimpse the various correlations that would otherwise be obscured. We think we have achieved this in this paper, but clearly an even more expanded (and expensive!) investigation would be better.
@@ 4 Explosion and Debris Morphology
The aspherical morphology of the explosion debris reflects the turbulence of the pre- and post-explosion dynamics that has been a central feature of CCSN theory for decades (Burrows et al., 1995)444The Standing Accretion Shock Instability (SASI) (Blondin et al., 2003), a vortical-acoustic instability (Foglizzo et al., 2007), once seemed manifest in most CCSN simulations. However, the associated up and down “sloshing” was in most models an artifact of the axial effect in 2D simulations; the turbulence witnessed was in fact neutrino-driven convection. However, Foglizzo et al. (2006) has shown that when the stalled shock recedes a SASI can emerge, but current 3D modeling (Burrows et al., 2020, 2023a) reveals it likely to be the spiral SASI (Blondin & Shaw, 2007) (However, see Melson et al. (2015) for an exception). Its emergence can presage a late-time explosion (Andresen et al., 2019), an explosion followed by the formation of a black hole (Burrows et al., 2023a), or (as we see quite often (Burrows et al., 2023a)), a fizzled collapse that will over time lead to the “quiescent” collapse to a black hole.. Bubble structures are universally in evidence (see Sato et al. (2021)), with larger bubbles generally driving the explosions. Importantly, the longer the delay to explosion, the more vigorous the turbulence achieved, with the Mach number of the convective overturning motions prior to explosion getting as high as 0.5. However, if the explosion is more prompt, the post-shock turbulence does not have sufficient time to grow to such vigor and large convective structures are not produced. The result in such cases is a more spherical explosion, albeit with many small bubble structures. Given the trends of compactness and progenitor ZAMS mass with explosion time exposed in Figure 5, we would expect lower compactness, lower ZAMS mass progenitors would explode more spherically, while those with larger compactness would explode more aspherically.
@@
Such structures have a direct bearing on the relative line profiles and the spatial distribution of the elements in supernova debris (Taubenberger et al., 2009; Sato et al., 2021; Fang et al., 2023) and the associated (anti)correlation of Ni and progenitor oxygen shell material is a qualitative prediction of these 3D models. Moreover, one would expect a correlation with the polarization of the inner debris (Leonard et al., 2000, 2006; Tanaka et al., 2017; Nagao et al., 2023). Figure 9 portrays the relationship between the explosion energy and the dipole anisotropies of the ejecta mass (left) and the ejecta energies (right) among our set of twenty models. The dipoles are the amplitudes of the spherical harmonic components of their angular distributions. These panels suggest a roughly monotonic correlation of two indices of anisotropy with explosion energy that is a new prediction that emerges from our study. We expect that, though with the common caveat of some scatter, more energetic explosions to evince greater anisotropy, but this needs to be tested.
@@ 5.2 Explosion Energy
Two of the most important observables of CCSN theory are the explosion energy and the gravitational mass of the residual neutron star. Our theoretical results reveal that the two are tightly coupled. Figure 11 portrays the relation between the two and indicates that, with the caveat that some of our models have not quite asymptoted to their final explosion energies (but are generally within 10% or better), the explosion energy is an almost monotonic function of neutron star gravitational mass. This relationship between two observables is one of the central results of this paper and speaks to the importance of our philosophy to explore many long-term 3D state-of-the-art simulations to extract meaningful predictions. This strong mapping between observables would not have been clearly discerned without a 3D investigation of many progenitors to late times and should inform future observational campaigns to extract physically useful parameters of core-collapse supernova explosions. The study of but one single model would not be so illuminating.
Figure 12 plots the connection between binding energy and explosion energy (left) and compactness and explosion energy (right). Both relationships are roughly monotonic and robust, but the explosion energy/binding energy correlation is physically one of the most important findings from the theoretical perspective (Janka, 2017a). It implies that the supernova energy is a “self-regulating” phenomenon, akin in broad outline to winds. Such self-regulation, mediated through the shared implicit dependence upon compactness, may become one of the unifying principles of CCSN theory going forward.
@@ 5.3 Ni Yields and Nucleosynthesis
Supernova nucleosynthesis is a vast subject that spans almost eight decades. Components are pre-supernova progenitor yield calculations, explosive nucleosynthesis, -rich freeze-out, and various sub-processes of differing import. Reviews can be found in Woosley et al. (2002); Nomoto et al. (2013); Fischer et al. (2023); Arcones & Thielemann (2023). The decades-long tradition of 1D models Thielemann et al. (1996); Rauscher et al. (2002); Woosley et al. (2002); Heger & Woosley (2010); Nomoto et al. (2013); Sukhbold et al. (2016); Limongi & Chieffi (2018); Curtis et al. (2019) is now being challenged by 3D models (Lentz et al., 2015; Wongwathanarat et al., 2015; Harris et al., 2017; Wanajo et al., 2018; Sieverding et al., 2020, 2023; Wang & Burrows, 2023a, b). Winds play an important role (Duncan et al., 1986; Burrows, 1987a; Burrows et al., 1995; Qian & Woosley, 1996; Wang & Burrows, 2023a), as do neutrinos (Woosley et al., 1990; Fröhlich et al., 2006; Pruet et al., 2006) and -ray emissions of the ejecta can play a diagnostic role (Diehl & Timmes, 1998; Diehl et al., 2021; Diehl & Prantzos, 2023).
One of the central observables of core-collapse supernova explosions is the Ni yield (in solar masses), since the Ni decay chain to Fe powers some of the supernova light curve and can easily be measured. Hence, the predicted Ni yields are central physical quantities of the CCSN phenomenon. Figure 13 provides the relation that emerges in our 3D simulation suite between explosion energy (in Bethes 10 ergs) and Ni production (in ). Ni production, and in general the ejecta isotopic abundances, are calculated using SkyNet (Lippuner & Roberts, 2017) with a 1530-isotope network including elements up to , together with about 340,000 post-processed backwardly-integrated tracers in each simulation. More details concerning the formalism can be found in Wang & Burrows (2023b). This too is one of the central results of this paper, but the correlation between the two has already been anticipated theoretically (e.g. Sukhbold et al. (2016)) and observed astronomically (Rodríguez et al., 2023). The left hand side provides the Ni correlation with explosion energy, with the range of predicted energies due to the fact that not all 3D models have yet asymptoted to their final values indicated. The right hand side shows the same numerical data, but superposed on the Rodríguez et al. (2023) measurements. The overlap between the two suggests a reasonable concordance, though both trends are subject to natural variation. Note that the strong theoretical trend of Ni mass with explosion energy includes the exploding black hole formers (triangles).
@@
Figure 16 shows the dependence of the freeze-out mass555The NSE temperature is set at 7 GK, and the freeze-out mass represents the total amount of matter that has ever achieved this temperature (Wang & Burrows, 2023b). upon compactness (top left), ejecta mass dipole (top right), and explosion energy (bottom). We see here too the rough monotonicity, in particular versus the explosion energy, of the mass in the freeze-out component. This follows the corresponding correlation of the explosion energy with Ni mass depicted in Figure 13 and is not surprising. It is nevertheless a strong prediction of the theory of CCSN explosions, as we see it, and is testable.